### v1

@@ Abstract
The Forward-Forward (FF) algorithm is a recent, purely forward-mode learning method, that updates weights locally and layer-wise and supports supervised as well as unsupervised learning. These features make it ideal for applications such as brain-inspired learning, low-power hardware neural networks, and distributed learning in large models. However, while FF has shown promise on written digit recognition tasks, its performance on natural images and time-series remains a challenge. A key limitation is the need to generate high-quality negative examples for contrastive learning, especially in unsupervised tasks, where versatile solutions are currently lacking. To address this, we introduce the Self-Contrastive Forward-Forward (SCFF) method, inspired by self-supervised contrastive learning. SCFF generates positive and negative examples applicable across different datasets, surpassing existing local forward algorithms for unsupervised classification accuracy on MNIST (MLP: 98.7%), CIFAR-10 (CNN: 80.75%), and STL-10 (CNN: 77.3%). Additionally, SCFF is the first to enable FF training of recurrent neural networks, opening the door to more complex tasks and continuous-time video and text processing.
@@
Local learning, Self-supervised learning, Unsupervised learning, Contrastive learning, Neuromorphic Computing
@@ 1 Introduction
Purely-forward and local algorithms propagate signals from inputs to outputs, without any backward pass, and employ a learning rule that is local in space, using signals only from pre- and post- neurons to update a weight. These algorithms, which include well-known Hebbian learning methods [1], have long been studied to model the brain, as their features align with biological processes, contrarily to error backpropagation [2]. They are also ideal for on-chip learning of low-power embedded hardware. The purely forward architecture addresses the challenge of implementing backpropagation on neuromorphic chips, a key obstacle that has been limiting the development of on-chip training for these systems [3]. The local learning rule allows synaptic devices to be directly updated by signals from neuron devices, eliminating the need to store neural activations for gradient computation, thus significantly reducing memory usage, power consumption as well as training times [4].
The practical application of these algorithms was historically limited by low accuracy on deep learning image recognition benchmarks, such as CIFAR-10 (labeled) and STL-10 (unlabeled). However, in recent years, accuracy has significantly improved thanks to their integration with automatic differentiation tools and activation functions, combined with the use of standardization and pruning techniques [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. As a result, these algorithms have become relevant not only for brain modeling and energy-efficient training of neuromorphic chips, but also for distributed learning of large-scale models across multiple devices [15, 16, 17, 18].
These advances have motivated the recent development of novel purely forward and local algorithms aimed at improving accuracy and simplifying implementation for various applications. The Forward-Forward (FF) algorithm is a key example of this new breed of algorithms [19]. Its main advantages are its simplicity and versatility, allowing it in principle to handle both unsupervised and supervised learning, and to process time-series as well as static inputs — while many other purely forward and local algorithms are limited to a single category. As a result, the FF algorithm has attracted significant interest since its introduction [20, 21, 22, 7, 23, 24, 25, 26, 27], inspiring variants with improved accuracy and architectures and addressing tasks such as image recognition, temporal sequence classification, and reinforcement learning [28, 29, 30, 14, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Early demonstrations of FF in silico already cover a wide range of hardware platforms, including microcontrollers [40], in-memory computing systems utilizing Vertical NAND Flash Memory as synapses [41], and physical implementations where neurons are replaced by acoustic, optical, or microwave technologies [42, 43].
Figure 1: Comparison of algorithms that are both purely forward and local across different tasks. The blue frame displays the CIFAR-10 results achieved by various supervised local learning methods. The green frame presents the results of different local and unsupervised learning methods on CIFAR-10, STL-10, and MNIST datasets, respectively. Algorithms related to FF are highlighted in purple. FF1 refers to [19], FF2 to [33], FF3 to [14], FF4 to [44], and FF5 to [38]. Our method, SCFF, is indicated in red. SigProp refers to [5], DRTP to [6], Act.Learning to [7], PEPITA to [8], SoftHebb to [9], CSNNs to [10], HardHebb2 to [12], HardHebb1 to [11], Hebb+WTA to [9] and Hebb-CHU to [13].
Fig 1 highlights recent advancements in the Forward-Forward algorithm (FF, shown in purple) [19, 33, 14, 38], comparing its accuracy on image classification tasks with algorithms that are also both purely forward and local (shown in black [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]). As illustrated, the accuracy of FF supervised training has significantly improved from below 60% [19] to 88.2% recently [44], progressively catching up with the best supervised, purely local, and forward algorithm [5]. In unsupervised learning, where FF’s greatest potential for on-chip training may lie, it has only achieved high accuracy on the MNIST task [45, 19], with limited [38] or no success on more challenging datasets such as CIFAR-10 [46] or STL-10 [47].
Inspired by the noise contrastive estimation (NCE) method [48], the Forward-Forward (FF) algorithm presents positive and negative examples sequentially to the network inputs. Once trained, the network is expected to produce distinctly different neural responses across all layers for these examples. The key challenge is generating“negative” examples that closely resemble the training data but still provide enough contrast for the network to learn meaningful representations.
Supervised learning methods applicable to any dataset have been developed, solving tasks like MNIST and CIFAR-10 [19, 32]. However, for unsupervised FF learning, there is no universal method to generate positive and negative examples for all databases, hindering FF’s application to more complex unsupervised tasks beyond MNIST. This limitation is evident in the few FF points in the Unsupervised panel of Fig 1, where only Hebbian and activation learning algorithms have successfully solved CIFAR-10 and STL-10 by combining a local rule with a purely forward architecture.
Moreover, the FF algorithm currently lacks the ability to handle time-varying sequential data, limiting its applicability in neuromorphic systems which often deal with dynamic inputs from the real world. While the original FF paper [19] includes a multi-layer recurrent neural network, its purpose is to model top-down effects, using a static MNIST image repeated over time frames as input. Another implementation demonstrates a limited form of sequence learning with a fully connected network, but this architecture could not handle real-time sequential data due to the absence of recurrence. As a result, FF has yet to be extended to effectively handle recurrent network scenarios for time-varying inputs.
In this work, we introduce the Self-Contrastive Forward-Forward (SCFF) method, where each data sample is contrasted within itself to enable efficient learning. This method is inspired by self-supervised learning to provide a simple and effective way to generate positive and negative examples for any dataset. In SCFF, a positive example is created by concatenating an input with itself, while a negative example is formed by concatenating the input with a randomly selected example from the training set. This simple method (red points in Fig 1) not only extends the capabilities of FF to unsupervised tasks but also surpasses the state-of-the-art accuracy of similar algorithms on the MNIST, CIFAR-10, and STL-10 image datasets. It also opens the path to solving sequential tasks by contrasting inputs with FF. More specifically, our contributions are:
@@
  • We propose an approach called SCFF that takes inspiration from self-supervised learning to generate positive and negative examples and train neural networks with the Forward-Forward algorithm in an unsupervised way, applicable to a wide range of databases.
@@
  • We show that SCFF significantly outperforms existing unsupervised purely-Forward and local learning algorithms in image classification tasks. With a multilayer perceptron (MLP), we achieved a test accuracy of 98.7% on MNIST. With a convolutional neural network (CNN), we reached 80.75% on CIFAR-10 and 77.3% STL-10 (which includes a small number of labeled examples alongside a much larger set of unlabeled data).
@@
  • We present the first demonstration of the FF approach being successfully applied to sequential data. Our findings show that the proposed SCFF method effectively learns representations from time-varying sequential data using a recurrent neural network. In the Free Spoken Digit Dataset (FSDD), SCFF training results in an accuracy improvement of over 10 percentage points compared to scenarios where hidden connections remain untrained (random).
@@
  • We conduct a theoretical and illustrative analysis of the distribution of negative examples generated with our method in comparison to positive examples within the data space. The analysis reveals that negative data points consistently position themselves between distinct clusters of positive examples. This positioning suggests that negative examples play a crucial role in pushing apart and further separating adjacent positive clusters, thereby enhancing the efficiency of classification.
@@
Our results demonstrate the potential of Self-Contrastive Forward-Forward (SCFF) methods for efficient and flexible layer-wise learning of useful representations in a local and purely forward unsupervised manner. In section 2.1 and 2.2, we will introduce the two foundational methods that SCFF builds upon: the original Forward-Forward algorithm and Contrastive Self-Supervised learning, highlighting their differences and similarities. Next, in section 3, we will present our Contrastive Forward-Forward algorithm and discuss our findings. Finally, we will explore the relationship of SCFF to other purely forward and/or local learning methods in the discussion of section 4.
2 Background
@@
Figure 2: Comparative diagram illustrating three distinct unsupervised (self-supervised) learning paradigms. a. Generation of a negative example is implemented by hybridization of two different images in the original FF paper [19]. b. In Forward Forward (FF) Learning, the layer-wise loss function is defined so as to maximize the goodness for positive inputs (real images) and minimize the goodness for negative inputs, each of which is generated by corrupting the real image to form a fake image, as shown in a. c. In Contrastive Learning, the InfoNCE loss function determines the similarity between representations of two inputs (two different inputs or two same inputs but with different augmentations) in the end of the network [49]. d. Our proposed Contrastive Forward Forward Learning algorithm combines the principles of Forward Forward Learning and Contrastive Learning algorithms to maximize the goodness for concatenated similar pairs and minimize the goodness for dissimilar pairs with a layer-wise loss funcion.
2.1 The original Forward-Forward algorithm
The original Forward-Forward algorithm is depicted in Fig 2b. For an input example $\boldsymbol{x}_{i}$, each layer’s output is assigned a ‘goodness’ score $G_{i}^{(l)}$, where $l$ is the layer index. This score is calculated as the mean of the squared activities of the output neurons at layer $l$: $G_{i}^{(l)}=\frac{1}{M^{(l)}}\sum_{m}{y_{i,m}^{2(l)}}$, where $M^{(l)}$ is the number of neurons at layer $l$ and $m$ represents the neuron index.

### v2

@@ Abstract
Agents that operate autonomously benefit from lifelong learning capabilities. However, compatible training algorithms must comply with the decentralized nature of these systems which imposes constraints on both the parameters counts and the computational resources. The Forward-Forward (FF) algorithm is one of these. FF relies only on feedforward operations, the same used for inference, for optimizing layer-wise objectives. This purely forward approach eliminates the need for transpose operations required in traditional backpropagation. Despite its potential, FF has failed to reach state-of-the-art performance on most standard benchmark tasks, in part due to unreliable negative data generation methods for unsupervised learning.
In this work, we propose Self-Contrastive Forward-Forward (SCFF) algorithm, a competitive training method aimed at closing this performance gap. Inspired by standard self-supervised contrastive learning for vision tasks, SCFF generates positive and negative inputs applicable across various datasets. The method demonstrates superior performance compared to existing unsupervised local learning algorithms on several benchmark datasets, including MNIST, CIFAR-10, STL-10 and Tiny ImageNet. We extends FF’s application to training recurrent neural networks, expanding its utility to sequential data tasks. These findings pave the way for high-accuracy, real-time learning on resource-constrained edge devices.
@@
On-chip learning, Local learning, Self-supervised learning, Unsupervised learning, Neuromorphic Computing
@@ 1 Introduction
On-chip edge learning is essential for adapting model weights in response to newly arrived, real-time, unlabeled personal data, making it especially suited for privacy-sensitive applications [1, 2]. Unlike centralized cloud training, edge training also significantly reduces energy consumption, making it a more efficient and sustainable choice [3, 4].
The surge in progress for inference using in-memory computing systems has yet to be matched in the realm of on-chip learning. Backpropagation [5], the highest-accuracy algorithm for training neural networks, is particularly challenging to implement in hardware. Its non-locality requires storing all activation functions and their derivatives, and the backward pass necessitates bidirectional crossbar arrays, which doubles the number of transistors or the number of arrays, resulting in higher power and area consumption [6]. Local learning algorithms, which compute weight updates locally within each layer or neuron without relying on dependencies across the entire network, offer a promising solution to these challenges. Thus, there is a pressing need for local algorithms that are not only efficient for hardware implementations but can also handle time-series data and perform unsupervised learning, all while maintaining high accuracy [7].
The upper section of Table 1 compares different categories of algorithms designed to address these challenges through their local updates, highlighting their specific advantages and issues for on-chip learning. While many algorithms effectively tackle specific aspects of on-chip learning, none achieve a comprehensive solution that simultaneously resolves all identified challenges.
Table 1: Comparisons of the learning capabilities of different local learning methods for online training in crosssbar based hardware and their test accuracy [%] on CIFAR-10, STL-10 and Tiny ImageNet datasets. For Tiny ImageNet, both the top-1 and top-5 accuracy are shown. The symbols , , and - mean “yes”, “no”, and “no reported results” respectively. ”Unidirectional inference” and ”Unidirectional learning” means that weight matrices do not need to be transposed during inference or learning. ”Unsupervised” means that the algorithm can handle unlabeled data. ”Time series” refers to the ability of algorithlms to handle time series input with RNN architecture.
Algori -thms Local Unidirec -tional inference Unidirec -tional learning Simple Activation function Unsuper -vised Time series CIFAR -10 STL -10 Tiny ImageNet
Top-1 Top-5
Backprop (backward pass) (ReLU) 93.8 [8] 91.70 [8] 54.8 [9]
Energy -based (bidirectional connections) (HardSigmoid) 92.3 [10] - -
Hebb -based (Oja’s weight transpose) (Softmax) 80.3 [11] 76.2 [11] - 37.0 [12]
Reward Hebb (reward backward) (ReLU) - 73.6 [13] -
DFA (random backward) (Tanh) 73.1 [14] - 32.1 [15]
FF (ReLU) 59.0 [16] - -
59.1 [17] - -
78.1 [18] - -
84.6 [19] - -
60.6 [20] - -
SCFF (ReLU) 80.8 77.3 35.7 59.8
Direct Feedback alignment (DFA) offers an alternative to backpropagation by simplifying the learning process. Instead of computing exact gradients, it uses fixed, random feedback weights [14, 21]. This approach has been widely adopted in various hardware implementations for on-chip learning [22, 23, 24, 25]. However, DFA still requires a backward pass through one or several synaptic arrays. Its performance also remains limited in deeper networks, particularly those integrating convolutional layers, despite impressive progress on other fronts [26].
Hebbian-based learning methods [27] present a strong interest for hardware implementations due to their simple two-factor learning rule that only depends on pre- and post- neuron activities. The recent development of algorithms such as SoftHebb [11] has considerably improved the accuracy of these approaches on unsupervised tasks. This success comes however at the cost of complex-to-implement softmax activation functions. Furthermore, bidirectional crossbar arrays are still needed for learning, as the weight update rule involves the transpose of the weight matrix.
Reward-based Hebbian algorithms incorporate a third factor in the Hebbian learning rule in order to increase the accuracy [13, 28, 29]. The reward signal flows from the output to the intermediate layers through additional weight matrices, thus complicating the implementation.
Energy-based algorithms [30, 31, 10, 32], such as Equilibrium Propagation [30, 33], offer spatially local updates, simple activation functions and accuracies very close to BP. Those models have been implemented in specialized hardware, using architectures like network of variable resistors [34] and Ising machines [35] to achieve efficient local updates and lower power consumption. However, they require hard-to-implement bidirectional connectivity as well as reaching network equilibrium for each input, which constitutes a hurdle for time-series classification.
Therefore, further work is essential to develop learning algorithms that integrate all the key features necessary for practical implementation: local learning rules, unidirectional inference and learning, simple activation functions, and high accuracy in embedded application tasks, such as unsupervised learning of both static and dynamic data.
Recently, the Forward-Forward (FF) algorithm [16] emerged as a promising candidate for hardware-friendly learning due to its simplicity: it is local, unidirectional, works with simple activation functions like ReLU, and supports both static and sequential data. As a result, the FF algorithm has attracted significant interest since its introduction [36, 37, 38, 39, 40, 17, 18, 41, 42, 43, 44, 20, 45]. Early demonstrations of FF in silico already cover a wide range of hardware platforms, including microcontrollers [46], in-memory computing systems utilizing Vertical NAND Flash Memory as synapses [47], and physical implementations where neurons are replaced by acoustic, optical, or microwave technologies [48, 49]. However, its unsupervised learning performance has stagnated, particularly due to the challenge of generating high-quality negative data.
The lower section of Table 1 highlights recent advancements in the Forward-Forward (FF) algorithm [16, 18, 39, 20, 17, 19], comparing its learning capabilities across various metrics. Notably, the accuracy of FF supervised training on CIFAR-10 dataset has improved significantly, rising from below 60% [16] to 84.7% [19] by integrating random direct feedback connections, thereby closing the gap with leading supervised, purely local, forward algorithms [50].
However, the performance of FF has plateaued in areas that are crucial for on-chip learning, such as unsupervised learning and time-series processing. In unsupervised learning, FF has shown only limited success [20] and has struggled with datasets more challenging than MNIST [51], such as CIFAR-10 [52], simplified ImageNet versions [53] or the unlabeled dataset STL-10 [54]. Additionally, FF is currently unable to effectively handle time-varying sequential data, which significantly limits its applicability to neuromorphic systems that often require processing dynamic, real-world inputs.
These challenges stem from the difficulty of generating “negative” examples that are similar enough to the “positive” training data to provide meaningful contrast, yet distinct enough for the network to learn effective representations. While creating such examples is relatively straightforward in supervised learning, where true or false labels can be integrated into the training process, there is currently no efficient, universal approach for generating positive and negative examples across all datasets. This limitation hinders FF’s effectiveness in unsupervised learning and time-series processing, making it difficult to generalize across different data types and applications.
In this work, we propose the Self-Contrastive Forward-Forward (SCFF) method, which contrasts each data sample against itself to enable efficient learning. Inspired by self-supervised learning, SCFF generates positive and negative examples directly from the data, making FF applicable to unsupervised learning across a wide range of datasets. We demonstrate that SCFF not only extends FF’s capabilities to unsupervised tasks but also surpasses other local algorithms for unsupervised tasks on the MNIST, CIFAR-10, STL-10 and Tiny Imagenet datasets. Moreover, SCFF makes it possible to apply FF to sequential tasks, unlocking its potential for time-series data processing. Specifically, our contributions are as follows:
@@
  • We propose the SCFF method, which efficiently generates positive and negative examples to train neural networks using the Forward-Forward algorithm in an unsupervised manner, applicable across a wide range of datasets.
@@
  • We show that SCFF significantly outperforms existing unsupervised local learning algorithms in image classification tasks, achieving test accuracies of 98.70% $\pm$ 0.01% on MNIST, 80.75% $\pm$ 0.12% on CIFAR-10, and 77.30% $\pm$ 0.12% on STL-10 (which includes a small number of labeled examples alongside a much larger set of unlabeled data). SCFF furthermore achieves a top-1 accuracy of 35.67% $\pm$ 0.42% and a top-5 accuracy of 59.75% $\pm$ 0.18% on the more complex dataset Tiny ImageNet.
@@
  • We present the first demonstration of the FF approach being successfully applied to sequential data. Our findings show that the proposed SCFF method effectively learns representations from time-varying sequential data using a recurrent neural network. In the Free Spoken Digit Dataset (FSDD), SCFF training improves accuracy by approximately 10 percentage points compared to reservoir computing methods.
@@
  • We conduct a theoretical and illustrative analysis of the distribution of negative examples within the data space. The analysis reveals that negative data points consistently position themselves between distinct clusters of positive examples, which suggests that negative examples play a crucial role in pushing apart and further separating adjacent positive clusters, thereby enhancing the efficiency of classification.
@@
Our results demonstrate the potential of Self-Contrastive Forward-Forward (SCFF) for efficient, layer-wise learning of meaningful representations in a local, purely forward, and unsupervised manner, enabling online training in hardware environments. In Section 2, we introduce SCFF, detailing its hardware applicability, innovative negative example generation method and its training procedure. Next, in section 3, we will present the results and discuss our findings. Finally, we will explore the relationship of SCFF to other purely forward and/or local learning methods in the discussion of section 4.
@@
Figure 1: Comparative diagram illustrating three distinct unsupervised (self-supervised) learning paradigms. a. Generation of a negative example is implemented by hybridization of two different images in the original FF paper [16]. b. In Forward Forward (FF) Learning, the layer-wise loss function is defined so as to maximize the goodness for positive inputs (real images) and minimize the goodness for negative inputs, each of which is generated by corrupting the real image to form a fake image, as shown in a. c. In Contrastive Learning, the InfoNCE loss function determines the similarity between representations of two inputs (two different inputs or two same inputs but with different augmentations) in the end of the network [55]. d. Our proposed Contrastive Forward Forward Learning algorithm combines the principles of Forward Forward Learning and Contrastive Learning algorithms to maximize the goodness for concatenated similar pairs and minimize the goodness for dissimilar pairs with a layer-wise loss function.
2 Self-Constrastive Forward-Forward algorithm
The primary innovation of SCFF lies in its method for generating positive and negative examples, making it applicable to both supervised and unsupervised tasks across a wide variety of datasets. SCFF maintains the hardware-friendly, forward-only nature of FF while significantly enhancing its learning capabilities, particularly for unsupervised tasks and time-series classification.
2.1 Creating the negative and positive examples
In the FF algorithm, illustrated in Fig. 1b, positive and negative examples are successively presented to the network input [16]. A “goodness” score, related to local neuron activity, is computed at each layer, with the training objective being to maximize goodness for positive examples while minimizing it for negative ones. Therefore, the negative examples must be carefully crafted to effectively challenge the network, requiring more sophisticated approaches than basic noise injection or occlusion. For image data, the negative examples should maintain similar short-range correlations to the original data but differ significantly in long-range correlations to produce distinct shapes and categories. For the MNIST dataset, Hinton proposed generating negative examples by applying masks to different digit images and combining them to create hybrid images that look like digits but are not digits (Fig. 1a). While this method works well for MNIST [16], it does not easily extend to more complex image databases like CIFAR-10, ImageNet and STL-10, or time series data, resulting in limited accuracy on these benchmarks. Therefore, further development of the FF algorithm must focus on devising robust methods for generating positive and negative examples that are applicable across diverse datasets.
As illustrated in Fig. 1c, contrastive self-supervised learning methods share some similarity with FF. Negative and positive pairs are defined, where a positive pair consists of two different augmented views of the same data sample, and a negative pair of two different samples. Contrastive losses are employed to define the similarity in feature space between each image of a pair, with the training objective being to maximize similarity for positive pairs while minimizing it for negative ones.
In SCFF, we propose to directly take pairs of positive and negative images as inputs to the neural network (Fig. 1d) instead of contrasting their representations in feature space as done in contrastive self-supervised learning (Fig. 1c). More specifically, given a batch of $N$ training examples, and for a randomly selected example $\boldsymbol{x}_{k}$ ($k\in\{1,N\}$) in the batch, the positive example $\boldsymbol{x}_{i,\text{pos}}^{(0)}$ (the number 0 is the layer index) is the concatenation of two repeated $\boldsymbol{x}_{k}$, i.e., $\boldsymbol{x}_{i,\text{pos}}^{(0)}=[\boldsymbol{x}_{k},\boldsymbol{x}_{k}]^{T}$. The negative example $\boldsymbol{x}_{j,\text{neg}}^{(0)}$ is obtained by concatenating $\boldsymbol{x}_{k}$ with another example $\boldsymbol{x}_{n}$ ($n\neq k$) in the batch, i.e., $\boldsymbol{x}_{j,\text{neg}}^{(0)}=[\boldsymbol{x}_{k},\boldsymbol{x}_{n}]^{T}$ (or $[\boldsymbol{x}_{n},\boldsymbol{x}_{k}]^{T}$ ). Fig. 2a shows some instances of generated positive and negative examples from the original training batch for the STL-10 dataset.
Importantly, although the input size is doubled, the computational cost and memory usage of network training are not impacted compared to standard FF because the weight matrices connecting $\boldsymbol{x}_{k}$ and $\boldsymbol{x}_{n}$ can be identical. For example, in the fully connected network illustrated in Fig. 1d, the outputs for the positive and negative examples from the first layer can be respectively written as:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{y}_{i,\text{pos}}^{(0)}=f(W_{1}\boldsymbol{x}_{k}+W_{2}\boldsymbol{x}_{k}),\quad\boldsymbol{y}_{j,\text{neg}}^{(0)}=f(W_{1}\boldsymbol{x}_{k}+W_{2}\boldsymbol{x}_{n}),}$ (1)
where $f$ is the ReLU activation function, and the weight matrices $W_{1}$ and $W_{2}$ correspond to the connections for each of the two images constituting the input. In practice, we set $W_{1}=W_{2}$ because the gradient of the loss function with respect to $W_{1}$ and $W_{2}$ converges to the same value. This results in a reformulation where the inputs are summed rather than concatenated, yielding:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{y}_{i,\text{pos}}^{(0)}=f(W(\boldsymbol{x}_{k}+\boldsymbol{x}_{k})),\quad\boldsymbol{y}_{j,\text{neg}}^{(0)}=f(W(\boldsymbol{x}_{k}+\boldsymbol{x}_{n})),}$ (2)
where $W=W_{1}=W_{2}$. This reformulation ensures that the effective input size remains unchanged compared to a standard single-image input, and the feature map dimensions in subsequent layers are not affected by the initial pairing operation. Unlike naive concatenation, which would increase feature map width, our approach maintains computational efficiency while preserving the relational structure of the input pairs.
Intuitively, this can be understood by recognizing that swapping the positions of $\boldsymbol{x}_{k}$ and $\boldsymbol{x}_{n}$ in the input should not affect the output neural activities. A rigorous mathematical proof of the convergence of matrices $W_{1}$ and $W_{2}$ is provided in Appendix A.
2.2 Training procedure
The training procedure builds upon the FF framework, enhanced by additional unsupervised training techniques:
  • Greedy Layer-wise Training / Joint Training: SCFF supports two training paradigms: greedy layer-wise training and joint training, allowing for flexibility in optimization strategy based on computational constraints and architectural requirements. In greedy layer-wise training, each layer of the network is fully trained before proceeding to the next layer. In joint training, all layers are updated simultaneously using SCFF’s local learning rule in each iteration. After unsupervised training with SCFF, we froze the network and trained a linear downstream classifier [56, 11] with the back-propagation method on representations created by the network using the labeled data. The linear classifier was optimized using cross-entropy loss. The accuracy of this classification serves as a measure of the quality of the learned representations.
  • Goodness Function: SCFF employs a ‘goodness’ function at each layer, similar to FF. The goodness score, $G_{i}^{(l)}=\frac{1}{M^{(l)}}\sum_{m}{y_{i,m}^{2(l)}}$, where $l$ is the layer index and $m$ represents the neuron index, calculated as the sum of squared activations, is optimized such that positive examples have higher goodness than negative examples.